# JudgeAgent: Knowledge-wise and Dynamic LLM Evaluation with Agent-as-Interviewer

### v3

@@ 3.3 Evaluation Feedback
This process constitutes the result generation phase. JudgeAgent aggregates the target’s all Q&A performances in the above two stages, and evaluates the target from multiple dimensions, including knowledge boundaries, logical reasoning, and comprehensive performance, producing a text evaluation report. Based on this evaluation, JudgeAgent simultaneously provides suggestions aimed at addressing the identified capability gaps, serving as a reference for model optimization.
Inspired by Generative Reward Model methods (Ankner et al., 2024; Ye et al., 2024; Li et al., 2025) and the evaluation report method proposed by (Kim et al., 2025), JudgeAgent generates feedback and suggestions based on all Q&A history for each batch, without directly providing the correct answer or background knowledge. Benefiting from context graphs, the suggestions can identify key knowledge concepts and extend relevant information to provide concise and valuable feedback, rather than providing ”cheating information” specific to seed questions. To verify the suggestion’s effectiveness, the target model is prompted to answer the same questions again with these suggestions. By comparing accuracy before and after the intervention, we can indirectly validate the effectiveness of the evaluations.
@@ 4 Experiments
In this section, we validate the effectiveness of JudgeAgent’s evaluations using the method in Section 3.3, which is detailed in Algorithm 2. The following research questions guide our experiments: First, does JudgeAgent genuinely discover the shortcomings of the target model (RQ1)? Second, to what extent does each mechanism in JudgeAgent influence the evaluations (RQ2)?
Experiment Setup. We select GPT-4.1111We use gpt-4.1-2025-04-14 from OpenAI’s official API for all experiments(Achiam et al., 2023) to be the core LLM in JudgeAgent, which serves as both a generator and evaluator model. In our experiments, the batch size in Benchmark Grading is 3, and the Interactive Extension is limited to a maximum of 3 rounds.
@@ 4.1 Dataset and Target Model Selection
The initial stage of JudgeAgent, Benchmark Grading, requires a predefined static benchmark. In our experiments, we select MedQA(Jin et al., 2021), MultiHop-RAG(Tang & Yang, 2024), and QuALITY(Pang et al., 2022). We remove the background knowledge of questions from MedQA and MultiHop-RAG during evaluation to evaluate the target’s knowledge boundaries rather than its comprehension ability, and to verify whether JudgeAgent can discover the target’s knowledge deficiencies. We use QuALITY to validate the effectiveness of JudgeAgent in guiding the comprehension and reasoning of target models. For detailed statistical information and the preprocessing procedures of these datasets in our experiments, refer to Appendix D.
For the target models, we select three LLMs as the targets: Qwen3(Yang et al., 2025), GLM4-Flash(GLM et al., 2024), GPT-4.1(Achiam et al., 2023), and gemini-2.5-pro(Comanici et al., 2025) [^2] We use qwen-plus-2025-04-28, gpt-4.1-2025-04-14, gemini-2.5-pro-preview-06-05, and the free version glm-4-flash-250414 as the target models.. We utilize the official APIs to interact with these models.
@@ 4.2 Evaluation Metrics
(1) Accuracy (ACC). We use this metric to measure the performance of the target model on base datasets, which is the proportion of questions correctly answered by the target model. To investigate the validity of JudgeAgent’s evaluations, we compared the changes in the ACC before and after dynamic evaluation. ACC1 represents the target model’s ACC in answering base questions in Benchmark Grading stage, measuring its performance before receiving evaluation feedback. ACC2 denotes the ACC in answering the same questions after receiving feedback from JudgeAgent.
(2) Correction Rate (CR). This metric quantifies the proportion of questions that the target model initially answered incorrectly but subsequently answered correctly after receiving evaluation suggestions, which measures the effectiveness of the feedback suggestions from JudgeAgent. A higher correction rate indicates better performance of the suggestions.
(3) Correct-to-Error Rate (CtE). This metric serves as the inverse of the Correction Rate, representing the proportion of questions that were initially answered correctly but subsequently answered incorrectly. A lower CtE indicates greater effectiveness of the suggestions.
@@ 4.3 Main Results and Analysis
Based on the results from the knowledge-intensive datasets, MedQA and MultiHopRAG, as presented in Table 2, JudgeAgent can effectively identify potential knowledge gaps in the target LLM and subsequently mitigate these gaps by providing targeted prompts that indicate possibly missing knowledge points or overlooked knowledge associations to the target models. Furthermore, as evidenced by the overall performance on QuALITY, a dataset emphasizing reasoning and comprehension, as presented in Table 2, JudgeAgent also contributes to providing further optimization guidance to address potential shortcomings in the logical reasoning and semantic understanding abilities of the target models, thereby assisting in refining the target’s thinking steps. Additionally, by comparing the Correction Rate (CR) and the Correction-to-Error Rate (CtE) of different models before and after receiving evaluation suggestions, it can be observed that the effectiveness of the suggestions is less consistent for relatively weaker models (e.g., the free model GLM4-Flash), as reflected in the higher CtE. In contrast, stronger models are less susceptible to misleading suggestions.
@@
Based on the performance across different difficulties of QuALITY as shown in Table 2, we can analyze the effectiveness of JudgeAgent’s suggestions for different targets on various difficulties. For stronger models (Qwen3, GPT-4.1, and Gemini-2.5-pro), JudgeAgent provides stronger guidance and optimization, particularly on questions of higher difficulty (Medium and Hard), since these models may have already mastered the basic knowledge of Easy-level questions. In contrast, for the weaker models, JudgeAgent leads to considerable improvement across all difficulty levels, with particularly notable gains on the Easy level, since JudgeAgent’s feedback is more effective at filling gaps in basic concepts. These results further indicate that JudgeAgent more accurately assesses the capability boundaries of targets and provides more difficulty-adaptive optimization to the targets. Furthermore, for all target models, JudgeAgent’s suggestions demonstrate clear optimizing guidance on Hard-level questions that involve complex reasoning, suggesting that JudgeAgent effectively identifies underlying deficiencies in target models through dynamic interactive evaluation.
In summary, the results indicate that JudgeAgent can effectively identify potential knowledge or capability gaps in target models and optimize their performance by providing targeted suggestions. Furthermore, through a difficulty-adaptive, dynamic interaction-based evaluation, JudgeAgent can more precisely delineate the target model’s capabilities, offering more refined evaluation results.
In addition, we conducted cross-validation experiments to verify that the JudgeAgent’s suggestions are effective not only for base questions, further corroborating the validity of the main experiment results. The results and analysis are detailed in Appendix E.2
@@ 4.4 Ablation Study
To address RQ2, we conduct ablation studies on MedQA with GLM4-Flash as the target LLM. When maintaining the responses to base questions unchanged, we removed different modules of JudgeAgnet. By comparing the performance improvement under various ablation settings, we investigated how much different modules influence the evaluations. The results are shown in Table 3.
Table 3: The results of the ablation study with MedQA as the base dataset, and all values are percentages. $\Delta=\text{ACC2}-\text{ACC1}=\text{CR}-\text{CtE}$ refers to the overall improvement on the target models after receiving evaluation from JudgeAgent.
Target Model Evaluator MedQA
[…]
GLM4-Flash JudgeAgent 80.09 92.82 13.46 0.73 12.73
   -w/o context graph 88.60 10.26 1.75 8.50
   -w/o difficulty-adaptive 89.68 11.14 1.56 9.59
   -w/o interactive extension 86.50 9.47 3.06 6.41
GPT-4.1 JudgeAgent 84.97 92.44 7.65 0.18 7.47
   -w/o context graph 91.36 6.72 0.33 6.39
   -w/o difficulty-adaptive 91.94 7.36 0.38 6.98
   -w/o interactive extension 89.81 6.13 1.28 4.85
Qwen3 JudgeAgent 91.71 96.38 5.02 0.35 4.67
   -w/o context graph 95.57 4.28 0.42 3.86
   -w/o difficulty-adaptive 95.94 4.59 0.36 4.23
   -w/o interactive extension 93.01 1.97 0.67 1.30
JudgeAgent (w/o context graph) removes the context graph, which ensures the knowledge relevance between synthesized questions and the base ones, and only uses chunks sampled randomly from the original knowledge base. The results indicate that removing the context graph reduces the improvement effect of the JudgeAgent’s evaluations. A possible explanation is that the lack of context graph prevents extended questions and base questions from being associated in terms of focused entities and knowledge, leading to fragmented suggestions without accurate information, which may disrupt the thinking of target LLMs. The higher correction-to-error rate (CtE) compared to the setting -w/o difficulty-adaptive also provides supporting evidence for this potential interference.
JudgeAgent (w/o difficulty-adaptive) removes the difficulty-adaptive mechanism, which enables JudgeAgent to assess the target model’s capability precisely, and generates questions with fixed difficulty rules in the prompt. The results show that the removal of this module diminishes the effectiveness of JudgeAgent’s evaluations, demonstrating the importance of the difficulty-adaptive mechanism for providing effective evaluations.
JudgeAgent (w/o interactive extension) removes the Interactive Extension, which is the core process of dynamic evaluation to extend the evaluation scope of knowledge and capability, and only evaluates models with base datasets. The results show that removing the Interactive Extension significantly weakens the effectiveness of JudgeAgent compared to other ablation settings. This indicates that the evaluation enabled by the dynamic expansion, which expands both breadth and depth of knowledge and capability in dynamic evaluation, is crucial to the final evaluation suggestions.
Additionally, the ablation experiment results comparing different models reveal that for relatively weaker models (e.g., GLM4-Flash), the performance decline after removing each component is more pronounced, which indicates that these dynamic evaluation mechanisms are particularly beneficial to weaker models and can effectively facilitate their optimization.
@@ 4.5 Parameter Analysis
Figure 3: The results of different expansion rounds on MedQA and MultiHop-RAG. @K indicates the ACC improvement after receiving suggestions from the K-th interaction.
Are more expansion rounds better for JudgeAgent? To answer this research question, we tested the evaluation effectiveness of JudgeAgent under different rounds in the Interactive Extension, as shown in Figure 3. The results demonstrate that as the expansion rounds increase, the effectiveness of JudgeAgent, which is represented by the accuracy improvement of target models, also gradually improves. However, the trend gradually slows down, showing a relatively clear marginal effect.
As the rounds increase, JudgeAgent can expand more questions around the knowledge related to base questions. After reaching certain rounds, the Q&A history is sufficient for JudgeAgent to identify the target’s knowledge deficiencies around base questions. Additional questions serve only as corroborative rather than critical evidence. There is a clear marginal effect in Figure 3. For example, the results of GPT-4.1 on MultiHop-RAG, the improvement per round decreases gradually from 2.04% by the first round, 1.13% by the second, 0.71% by the third, and to 0.23% by the last round, and other curves exhibit similar patterns. Therefore, JudgeAgent expands a maximum of 3 rounds in our experiments to avoid wasting resources and time.
@@ 4.6 Case Study
To gain a deeper understanding of Agent-as-Interviewer’s mechanism, we analyze GLM4-Flash’s responses to MedQA questions by receiving suggestions from direct evaluation and Agent-as-Interviewer. The seed questions where the target answers correctly are omitted in this case. A brief overview of the case is shown in Figure 4, while the detailed content is provided in Appendix H.
Figure 4: The brief overview of the comparative case in the Case Study.
In this case, the target LLM answers the base question incorrectly. When directly evaluated by the LLM evaluator, the evaluator adopted a more conservative approach to generically enumerate the potential consequences of closed abdominal injury, rather than to perform a comparison between the specific medical concepts involved. Consequently, even after receiving the feedback, the target LLM still outputs an incorrect answer, which is attributable to a nuance it failed to discern: peritonitis symptoms resulting from both colon injury and duodenal bulb injury can exhibit delayed onset.
In contrast, JudgeAgent first extracted two key entities, closed abdominal injury and peritonitis. It then retrieved relevant entities with texts, such as splenic rupture and duodenal injury, from the context graph. Based on the sampled knowledge paths, JudgeAgent generated a series of extended questions. Given that the target model consistently errs on all questions where duodenal injury and closed abdominal injury co-occur, JudgeAgent can determine that the target model had insufficient understanding of this concept entity, thereby providing specific knowledge guidance. Ultimately, the target LLM successfully answered the original base question correctly with this feedback.
The case above demonstrates that through the Agent-as-Interviewer paradigm, JudgeAgent is capable of accurately identifying gaps in the target’s knowledge and providing targeted feedback.
@@ 5 Conclusion
In this paper, we propose Agent-as-Interviewer, a dynamic evaluation paradigm that employs LLM agents to conduct multi-turn interactions for evaluation. This paradigm utilizes agents to call knowledge tools for wider and deeper knowledge in the dynamic question generation to achieve more complete evaluations, and leverages agents to plan query strategies for adjustment of the question difficulty to match the actual capabilities of target LLMs. Based on this paradigm, we develop JudgeAgent, a knowledge-wise dynamic evaluation framework to assess LLMs in multi-turn Q&A sessions and provide valuable suggestions that assist target LLMs in optimizing themselves. Thorough experiments validate that Agent-as-Interviewer can precisely assess the target’s knowledge boundaries. In our future work, we will further refine this novel evaluation paradigm and develop more reliable evaluation frameworks for LLMs.

### v4

@@ 3.3 Evaluation Feedback
Inspired by Generative Reward Model methods (Ankner et al., 2024; Ye et al., 2024; Li et al., 2025) and the evaluation report method in (Kim et al., 2025), JudgeAgent generates a text evaluation report based on all Q&A history for each batch. Benefiting from context graphs, the evaluation report can identify the knowledge deficiencies of target LLMs and provide concise and valuable feedback around key knowledge concepts.
@@ 4 Experiments
In this section, we validate the effectiveness of JudgeAgent using the method in Section 3.3, which is detailed in Algorithm 2. The following research questions guide our experiments: RQ1, does JudgeAgent’s knowledge-driven evaluation genuinely discover the shortcomings of the target model? RQ2, is JudgeAgent’s dynamic evaluation more effective and precise than the static benchmarking? RQ3, to what extent does each mechanism in JudgeAgent influence the evaluations?
Current LLM knowledge evaluation methods lack quantifiable and standardized validation approaches. Therefore, to validate the effectiveness of JudgeAgent, we use a similar approach as (Kim et al., 2025). We use the evaluation report produced by JudgeAgent as suggestions to prompt the target LLM to answer the same questions again. Then we compare the accuracy before and after the intervention to indirectly validate the effectiveness. To avoid cheating, we produce suggestions without directly providing the correct answer or background knowledge of seed questions to JudgeAgent.
Experiment Setup. We select GPT-4.1111We use gpt-4.1-2025-04-14 from OpenAI’s official API for all experiments to be the core LLM. In our experiments, the batch size in Benchmark Grading is 3, and the Interactive Extension is limited to a maximum of 3 rounds.
@@ 4.1 Dataset and Target Model Selection
In our experiments, we select MedQA(Jin et al., 2021), MultiHop-RAG(Tang and Yang, 2024), and QuALITY(Pang et al., 2022) as the initial benchmarks. We remove the background knowledge of questions from MedQA and MultiHop-RAG to evaluate the target’s knowledge rather than its comprehension ability, and to verify whether JudgeAgent can discover the target’s actual knowledge deficiencies. We use QuALITY to validate the effectiveness in guiding comprehension and reasoning. The detailed information and the preprocessing procedures are in Appendix D.
For target LLMs, we select Qwen3(Yang et al., 2025), GLM4-Flash(GLM et al., 2024), GPT-4.1(Achiam et al., 2023), and gemini-2.5-pro(Comanici et al., 2025) [^2] We use qwen-plus-2025-04-28, gpt-4.1-2025-04-14, gemini-2.5-pro-preview-06-05, and the free version glm-4-flash-250414 as the target models.. We utilize the official APIs to interact with these models.
@@ 4.2 Evaluation Metrics
(1) Accuracy (ACC). We use this metric to measure the performance of the target model on base datasets. To investigate the effectiveness of JudgeAgent, we compared the changes in the ACC before and after dynamic evaluation. ACC1 represents the target’s ACC in answering base questions in Benchmark Grading, measuring its performance before receiving evaluation feedback. ACC2 denotes the ACC in answering the same questions after receiving feedback from JudgeAgent.
(2) Correction Rate (CR). This metric quantifies the proportion of questions that the target model initially answered incorrectly but subsequently answered correctly after receiving evaluation suggestions, which measures the effectiveness of JudgeAgent. A higher correction rate indicates better performance of the suggestions.
(3) Correct-to-Error Rate (CtE). This metric serves as the inverse of the Correction Rate. A lower CtE indicates greater effectiveness.
@@ 4.3 Main Results and Analysis
Based on the results from MedQA and MultiHopRAG in Table 2, JudgeAgent can effectively identify potential knowledge deficiencies in target LLMs and subsequently mitigate these deficiencies by providing targeted suggestions to the target LLM. Furthermore, as evidenced by the overall performance on QuALITY in Table 2, JudgeAgent also contributes to providing further guidance to address potential shortcomings in the logical reasoning abilities of the target LLM, thereby assisting in refining the target’s thinking steps. Additionally, by comparing CR and CtE of different LLMs before and after receiving suggestions, it can be observed that the effectiveness of suggestions is less consistent for relatively weaker LLMs (e.g., the free model GLM4-Flash), as reflected in the higher CtE. In contrast, stronger LLMs are less susceptible to misleading suggestions.
@@
Based on the performance across different difficulties of QuALITY in Table 2, we can analyze the effectiveness of JudgeAgent for different targets on various difficulty levels. For stronger LLMs (Qwen3, GPT-4.1, and Gemini-2.5-pro), JudgeAgent provides stronger guidance, particularly on questions of greater difficulty (Medium and Hard), since these LLMs may have already mastered the basic knowledge of Easy questions. In contrast, for weaker LLMs, JudgeAgent leads to considerable improvement across all difficulty levels, with particularly notable gains on Easy level, since JudgeAgent’s feedback is more effective at filling gaps in basic concepts. These results further indicate that JudgeAgent more accurately assesses the capability and provides more difficulty-adaptive guidance to the target LLMs. Furthermore, for all target LLMs, JudgeAgent’s suggestions demonstrate clear guidance on Hard questions, indicating that JudgeAgent effectively identifies underlying deficiencies in target LLMs through dynamic interactive evaluation.
In summary, JudgeAgent can effectively and precisely identify potential knowledge deficiencies in target LLMs, offering more refined evaluation results. Furthermore, we conducted cross-validation experiments to verify that the JudgeAgent’s evaluations are effective not only for seed questions, which are detailed in Appendix E.5. Additionally, experiments on the efficiency and stability of JudgeAgent are detailed in Appendix E.1 and E.2.
4.4 Analysis of JudgeAgent’s Dynamic Scores
Benchmark Target Model ACC(%) Difficulty Score Difficulty Distribution(%)
base @1 @2 @3 easy medium hard
MedQA GLM4-Flash 80.09 1.201 1.230 1.217 1.192 20.86 12.64 66.50
GPT-4.1 84.97 1.275 1.388 1.395 1.401 14.61 9.02 76.38
Qwen3 91.71 1.376 1.472 1.496 1.487 8.01 8.58 83.42
Gemini-2.5-pro 91.04 1.366 1.484 1.494 1.505 9.41 7.76 82.84
MultiHopRAG GLM4-Flash 51.25 0.769 0.587 0.535 0.486 34.06 48.13 17.81
GPT-4.1 68.94 1.034 1.006 1.000 0.996 11.55 40.16 48.29
Qwen3 63.65 0.955 0.801 0.767 0.746 19.62 48.91 31.47
Gemini-2.5-pro 62.25 0.934 0.967 1.005 1.019 13.63 38.80 47.57
QuALITY GLM4-Flash 74.35 1.065 1.143 1.232 1.304 12.09 39.30 48.60
GPT-4.1 91.04 1.320 1.428 1.522 1.603 1.86 23.49 74.65
Qwen3 88.53 1.296 1.388 1.466 1.520 3.26 25.58 71.76
Gemini-2.5-pro 93.57 1.397 1.502 1.595 1.686 0.00 17.91 82.09
Table 3: The dynamic difficulty scores and difficulty distribution of questions generated by JudgeAgent. base means the score for base static benchmark questions, and @K means the scores for extended questions at the K-th iteration.
To address RQ2, we analyze the dynamic difficulty scores and the difficulty distribution of generated questions during JudgeAgent’s Interactive Expansion stage. The results are shown in Table 3.
As the results on MedQA show, there is no significant difference in the accuracy and scores between GPT-4.1 and GLM4-Flash on the static benchmark (base). However, through multi-round dynamic evaluation, it can be observed that the difficulty of questions queried to GPT-4.1 continuously increases, while that for GLM4-Flash decreases, indicating that GPT-4.1’s actual knowledge mastery exceeds the preset difficulty of the static benchmark, further suggesting that GPT-4.1’s knowledge mastery is far higher than that of GLM4-Flash.
Similarly, on MultiHopRAG, Qwen3 and Gemini-2.5-pro show comparable performance on the static benchmark. While in dynamic evaluations, Qwen3’s performance is significantly lower than that of GPT-4.1 and Gemini-2.5-pro, with the gap progressively widening over rounds. This gap is also reflected in the fact that the generated questions for Qwen3 are primarily in Medium level. Meanwhile, although GPT-4.1 performs noticeably better than Gemini-2.5-pro on the static benchmark, their performance was comparable in dynamic evaluations, as can be seen from their similar difficulty distributions in the questions generated. Moreover, the multi-turn interview mechanism makes the differentiation in evaluation more pronounced. In summary, compared to static benchmarking, JudgeAgent’s dynamic evaluation can more precisely distinguish the actual knowledge mastery of different LLMs. Furthermore, we validate JudgeAgent’s resistance to data contamination through dynamic evaluation, with experiments and analysis detailed in Appendix E.3.
@@ 4.5 Ablation Study
To address RQ3, we conduct ablation studies on MedQA with GLM4-Flash as the target LLM. When maintaining the responses to seed questions unchanged, we removed different modules of JudgeAgnet. By comparing the results under different ablation settings, we investigate how much different modules influence the evaluations The results are shown in Figure 3.
Figure 3: The results of the ablation study with MedQA as the base dataset, and all values are percentages.
Figure 4: The brief overview of the comparative case in the Case Study.
JudgeAgent (w/o context graph) removes the context graph and only uses chunks sampled randomly from the knowledge base. The results indicate that removing the context graph reduces the effectiveness of the JudgeAgent. The lack of context graph severs the knowledge link between extended questions and base ones, leading to feedback without accurate information, which may disrupt the thinking of target LLMs. The higher CtE compared to the setting -w/o difficulty-adaptive also provides supporting evidence for this potential interference.
JudgeAgent (w/o difficulty-adaptive) removes the difficulty-adaptive mechanism and generates questions with fixed difficulty rules in the prompt. The results show that the removal of this module diminishes the effectiveness of JudgeAgent’s evaluations, demonstrating the importance of the difficulty-adaptive mechanism for providing effective evaluations.
JudgeAgent (w/o interactive extension) removes the Interactive Extension and only evaluates models with base benchmarks. The results show that removing this stage significantly weakens the effectiveness of JudgeAgent compared to other ablation settings, indicating that the dynamic expansion, which expands both breadth and depth of knowledge, is crucial to the evaluation’s effectiveness.
@@ 4.6 Parameter Analysis
Figure 5: The results of different expansion rounds on MedQA and MultiHop-RAG. @K indicates the ACC improvement after the K-th interaction.
The number of extension rounds is crucial for balancing the efficiency and effectiveness of JudgeAgent. We test the JudgeAgent’s effectiveness under different rounds in Interactive Extension, as shown in Figure 5. The results demonstrate that as the rounds increase, the JudgeAgent’s effectiveness, which is represented by the accuracy improvement of target LLMs, also gradually improves. However, the trend gradually slows down, showing a relatively clear marginal effect.
As the rounds increase, JudgeAgent can expand more questions related to the knowledge around base questions. After reaching certain rounds, the Q&A history is sufficient to identify the target’s knowledge deficiencies around base questions. Additional questions serve only as corroborative rather than critical evidence. There is a clear marginal effect in Figure 5. For example, the results of GPT-4.1 on MultiHop-RAG, the improvement per round drops from 2.04% to 1.13%, then to 0.71%, and eventually to 0.23%, and other curves exhibit similar patterns. Therefore, JudgeAgent expands a maximum of 3 rounds in our experiments to avoid wasting resources and time.
Additionally, parameter analysis experiments regarding batch size are presented in Appendix E.6.
@@ 4.7 Case Study
To further understand JudgeAgent’s mechanism, we analyze GLM4-Flash’s responses to MedQA questions by receiving suggestions from direct evaluation and JudgeAgent. The seed questions where the target answers correctly are omitted in this case. A brief overview of the case is shown in Figure 4, and the detailed content is provided in Appendix H.
In this case, the target LLM answers the base question incorrectly. When directly evaluated, the evaluator LLM generically enumerated the potential consequences of closed abdominal injury. Consequently, even after receiving the feedback, the target LLM still outputs an incorrect answer.
In contrast, JudgeAgent first extracted two key entities, closed abdominal injury and peritonitis. Based on the sampled paths on the context graph, JudgeAgent generated extended questions, which help discover the lack of understanding of duodenal injury and closed abdominal injury. Ultimately, the target LLM successfully answered the original question correctly with this feedback.
@@ 5 Conclusion
In this paper, we propose JudgeAgent, a knowledge-driven and dynamic evaluation framework for LLMs. To address the challenge of limited knowledge coverage, JudgeAgent leverages LLM agents with context graphs to traverse knowledge structures and employs knowledge-driven synthesis for question generation. Furthermore, to mitigate data contamination and difficulty mismatch, JudgeAgent dynamically adjusts the capability estimation and generates difficulty-adaptive questions based on the target’s responses in multi-turn interviews. Extensive experiments validate the effectiveness of JudgeAgent and highlight the significant potential of this knowledge-driven and dynamic evaluation paradigm in enhancing the evaluation results. In our future work, we will further refine this dynamic evaluation paradigm and develop more effective and reliable evaluation tools.
Limitations
In this paper, we validate JudgeAgent’s effectiveness and its great potential for enhancing the dynamic evaluation paradigm of LLMs through extensive experiments. However, due to space constraints, we focus on proposing a knowledge-driven and dynamic evaluation framework, and do not delve deeply into methods for optimizing the evaluated LLMs based on JudgeAgent. In Section 4 Experiments, we only employ simple prompt injection optimization as a validation approach to verify the evaluation’s effectiveness. In future work, we plan to further explore the potential of this evaluation paradigm, conducting in-depth research on how to perform targeted training and optimization of the evaluated LLMs based on JudgeAgent’s evaluation results, thereby establishing an efficient and comprehensive iteration cycle for LLMs.
Ethical Considerations
In this study, all datasets used, including MedQA, MultiHop-RAG, and QuALITY, were sourced in compliance with relevant usage guidelines, ensuring no violation of privacy. No personally identifiable information was used, and no experiments were conducted that could raise privacy or security concerns. We are committed to maintaining transparency and integrity throughout the research process.